The distribution of the binomial random variable (X) has the following parameters p = 0.3 and n = 9. Determine standard deviation


1
Expert's answer
2022-04-11T15:51:37-0400

"\sigma =\sqrt{np(1-p) }=\\=\sqrt{9\cdot0.3\cdot(1-0.3)}=1.375."